russ ran at 7 mph. janet ran at 5 mph. how long will it take for them to be a half mile apart
Found 2 solutions by checkley77, drk:
Answer by checkley77(12844)

D=RT
.5=(7-5)T
.5=2T
T=.5/2
T=.25 HOURS OR 15 MINUTES.
PROOF:
7*.25=1.75 MILES
5*.25=1.25 MILES.
1.75-1.25=.5 MILES IS THE DIFFERENCE.

Answer by drk(1908)

This is an RTD question. Here is a table based on the above information
person . . . . . . . rate . . . . . . . .time . . . . . . . .. distance
russ . . . . . . . . . . 7 . . . . . . . . . .y . . . . . . . . . . . 7y
janet . . . . . . . . . 5. . . . . . . . . . y. . . . . . . . . . . 5y
total . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .(1/2)
FRom column 3, we get
7y + 5y = 1/2
12y = 1/2
y = 1/24
1/24 hrs or ~ 2.5 minutes